OpenStudy (anonymous):

solve. 15x^4-28x^2+5=0

OpenStudy (anonymous):

(x^2 - 25)(x^2 - 3) (x + 5)(x - 5)(x^2 - 3) x + 5 = 0, x= -5 x-5 =0, x=5 x^2 - 3 = 0, x^2 = 3, x = +/- √3